# Nearly-polynomial inverse theorem for the $U^{d}$ norm in degree $d+1$

### 17 Mar 2026 v1

@@ 2.2 Proof outline
We start by sketching the standard proof of $\operatorname{GI}_{d}(d)$. The argument is roughly as follows. For $f$ of degree $d$, the $d$-th derivative $\Delta^{d}f$ is a degree-$d$ multilinear form, which is biased by assumption. Call that bias $\delta$. Using a partition-versus-analytic rank bound, such as Theorem 2.1, one obtains a partition-rank decomposition for $\Delta^{d}f$ of length $\log(\frac{1}{\delta})^{\tilde{O}_{d}(1)}$. By plugging in the diagonal $(x,\ldots,x)$ and dividing by $d!$, we obtain a decomposition of $f_{d}$ (the degree-$d$ part of $f$), meaning $\operatorname{rk}(f_{d})\leq\log(\frac{1}{\delta})^{\tilde{O}_{d}(1)}$. Writing $f=\sum_{i}\alpha_{i}\beta_{i}$, a standard Fourier/pigeonhole-type argument gives a correlation for $f_{d}$ of at least $\delta^{\tilde{O}_{d}(1)}$ with some member of the span of the $\alpha_{i}$ and $\beta_{i}$—which is of degree at most d1d-1—and thus a correlation for $f$ itself with such a lower-degree polynomial.
The proof of Theorem 2, for homogeneous polynomials, follows this classical recipe, except it utilizes a new correlation lemma (Lemma 3.1) to obtain a correlation with a linear combination of only the lower-degree factors from each product $\alpha_{i}\beta_{i}$ in the decomposition, which is a polynomial of degree as low as $\deg(f)/2$.
@@
In Section 4, we introduce a new notion of rank for polynomials called $\operatorname{rk}^{*}$. The objective here was to find a notion of rank for which the following holds: whenever $\operatorname{rk}^{*}(f)$ is small, one can find a large correlation of $f$ with a polynomial of degree at most $\deg(f)-2$ (whereas for the usual notion of rank, one can only generally get $\deg(f)-1$). It turns out that the requirement needed is quite weak: if $f=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ is a $\operatorname{rk}$-decomposition, for it to be a $\operatorname{rk}^{*}$ decomposition, we require the degree-$1$ polynomials in the decomposition to have linearly independent linear components. This requirement guarantees that the zero-locus of these polynomials is fairly large, which is essential for us to obtain significant correlation.
We now explain how the $\operatorname{rk}^{*}$-decomposition of $f$ is obtained. Write $g=f_{d+1}$ and $h=f_{d}$ for the homogeneous parts of degree $d+1$ and $d$ respectively. Proposition 5.1 provides a simple formula for $\Delta^{d}f$ for any polynomial $f$ of degree $d+1$. Proposition 6.4 shows that decomposing $g$ by the standard polarization process goes smoothly, thanks to the chain rule for bias (Lemma 6.3), which tells us that $\Delta^{d}g$} is biased. The decomposition of $h$ is a more delicate issue, since the polarization $\overline{h}$ need not be biased. However, we prove that $\overline{h}$ is biased when adding an appropriate partial derivative $\partial_{c}\overline{g}$; moreover, the sum $\overline{h}+\partial_{c}\overline{g}$ is a $d$-linear form, hence another application of Theorem 2.1 gives a decomposition to this sum. To deal with the summand $\partial_{c}\overline{g}$ we added, we use a chain rule for $\operatorname{rk}^{*}$ (Proposition 4.5).
@@
We finally remark that we cannot utilize the arguments leading to the polynomial bound for $\operatorname{GI}_{d}(d)$, in Theorem 3.4, in order to obtain a similar bound for degree $d+1$, since our argument requires more control over the factors than is afforded by that Theorem.
@@ 3 Correlation lemma, the homogeneous and the degree-$d$ cases
For a tuple of functions $\mathbf{A}=(A_{1},\ldots,A_{m})$, their joint zero set is $\operatorname{Z}(\mathbf{A})=\{x\,|\,\forall i\colon A_{i}(x)=0\}$. For an event $E$, we abbreviate $\operatorname{\mathfrak{bias}}(g\,|\,E)=\operatorname*{\mathbb{E}}[\chi(g)\,|\,E]$.
@@
Let $g,A_{1},\ldots,A_{m}\colon\mathbb{F}^{n}\to\mathbb{F}$ be any functions, and put $\mathbf{A}=(A_{1},\ldots,A_{m})$. Then
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\cdot\operatorname{\mathfrak{bias}}(g\,|\,\operatorname{Z}(\mathbf{A})).$
@@
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(g(x)-\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
$=\operatorname*{\mathbb{E}}_{x}\chi(g(x))\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
$=\operatorname*{\mathbb{E}}_{x}\chi(g(x))\mathbf{1}_{\operatorname{Z}(\mathbf{A})}(x)$
$=|\mathbb{F}|^{-n}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(g(x))$
$=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\frac{1}{|\operatorname{Z}(\mathbf{A})|}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(g(x))$
$=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\operatorname{\mathfrak{bias}}(g\,|\,\operatorname{Z}(\mathbf{A})).\qed$
We will also need the following Theorem due to Warning (for a short proof see [Asg23]).
@@
We obtain the following correlation bound, where $\operatorname{cor}_{d}(g)=\max_{\deg(P)\leq d}\operatorname{cor}(g,P)$.
@@
If $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(g)$ for some $\mathbf{A}=(A_{1},\ldots,A_{m})\in\operatorname{Form}_{d}(\mathbb{F})^{m}$, then $\operatorname{cor}_{d}(g)\geq|\mathbb{F}|^{-dm}$.
@@
Apply Lemma 3.1 to obtain
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{bias}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $\geq\Big|\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(g-\sum_{i=1}^{m}c_{i}A_{i}\Big)\Big|$
$=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\cdot\operatorname{bias}(g\,|\,\operatorname{Z}(\mathbf{A}))=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\geq|\mathbb{F}|^{-dm},$
where the last inequality applies Theorem 3.2, using that $0\in\operatorname{Z}(A_{1},\ldots,A_{m})$ by homogeneity. Thus, there exists a form $P=\sum_{i=1}^{m}c_{i}A_{i}$ of degree at most $d$ such that $\operatorname{bias}(g-P)\geq|\mathbb{F}|^{-dm}$. ∎
@@
Let $f:\mathbb{F}^{n}\to\mathbb{F}$ be a homogeneous polynomial with $k:=\deg(f)<2d$, where $\mathbb{F}=\mathbb{F}_{q}$ with $\operatorname{ch}(\mathbb{F}_{q})>k$, and put $\operatorname{bias}(\Delta^{d}f)=\delta$. First, we observe that $\operatorname{bias}(\overline{f})\geq\delta^{2^{d}}$. Indeed, using the monotonicity of Gowers norms, $\lVert{g}\rVert_{U^{k}}\geq\lVert{g}\rVert_{U^{d}}$ for $k\geq d$, and using the fact that, by definition, $\operatorname{bias}(\Delta^{t}f)=\lVert{e(f)}\rVert_{U^{t}}^{2^{t}}$ for any $t$, we have that
@@
with $Q_{i},R_{i}$ multilinear forms of degrees at most k1k-1, and
$r\leq\tilde{O}_{k}(\log_{p}(1/\operatorname{bias}(\overline{f})))=\tilde{O}_{d}(\log_{q}(1/\delta)).$
Assume without loss of generality that $\deg(Q_{i})\leq\deg(R_{i})$ for every $i$. Since $\deg(Q_{i})+\deg(R_{i})=k$, we have $\deg(Q_{i})\leq k/2$. Thus, $\overline{f}\in\langle Q_{1},\ldots,Q_{r}\rangle$. Plug in the diagonal: let $A_{i}(x)=Q_{i}(x,\ldots,x)$, and observe that $f(x)=\frac{1}{k!}\overline{f}(x,\ldots,x)$, using the homogeneity of $f$ and $\operatorname{ch}(\mathbb{F})>k$. It follows that $f\in\langle A_{1},\ldots,A_{r}\rangle$ with each $A_{i}$ homogeneous of degree at most $k/2<d$, by assumption. We are now done by Corollary 3.3:
$\operatorname{cor}_{<d}(f)\geq q^{-dr}\geq p^{-\tilde{O}_{d}(\log_{q}(1/\delta))}=\delta^{\tilde{O}_{d}(1)}.\qed$
@@ 3.1 A remark on the case of degree-$d$ polynomials
Let $f\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(f)\leq d$ and $\operatorname{bias}(\Delta^{d}f)=\delta$, where $\mathbb{F}=\mathbb{F}_{p}$ and $d<p$. Recall that $\Delta^{d}f\colon(\mathbb{F}^{n})^{d}\to\mathbb{F}$ is a multilinear form of degree $d$, and let us write $\Delta^{d}f(x^{(1)},\ldots,x^{(d)})=\sum_{i=1}^{n}x^{(d)}_{i}T_{i}(x^{(1)},\ldots,x^{(d-1)})$. Put $\mathbf{T}=(T_{1},\ldots,T_{n})$, and consider the multilinear variety $V=\operatorname{Z}(\mathbf{T})\subseteq(\mathbb{F}^{n})^{d-1}$. Note that the density of $V$ is $\mathbb{P}(\mathbf{x}\in\operatorname{Z}(\mathbf{T}))=\operatorname{bias}(\Delta^{d}f)=\delta$. By Theorem 3.4, there are multilinear forms $\mathbf{S}=(S_{1},\ldots,S_{m})$, with each $S_{i}\colon(\mathbb{F}^{n})^{d-1}\to\mathbb{F}$ of degrees at most d1d-1, such that $\operatorname{Z}(\mathbf{S})\subseteq\operatorname{Z}(\mathbf{T})$ ($\subseteq\operatorname{Z}(\Delta^{d}f)$) and $m\leq O_{d}(\log_{p}(1/\delta)+1)$. We observe that for any nonzero $d$-linear form $T$, $\operatorname{AR}(T)\geq 2^{-d}$; this follows for example from Claim 3.3 in [Lov19]. Since we may assume $\Delta^{d}f\neq 0$, as otherwise $\delta=\operatorname{bias}(\Delta^{d}f)=1$ and there is nothing to prove, we have that $\log_{p}(1/\delta)=\operatorname{AR}(\Delta^{d}f)\geq 2^{-d}$. Thus, $\log_{p}(1/\delta)+1\leq(2^{d}+1)\log_{p}(1/\delta)$, and therefore
@@
Plugging in the diagonal, let $A_{i}(x)=S_{i}(x,\ldots,x)$, homogeneous with $\deg(A_{i})\leq\det(S_{i})\leq d-1$. Put $\mathbf{A}=(A_{1},\ldots,A_{m})$. We have $f_{d}(x)=(1/d!)\Delta^{d}f(x,\ldots,x)$ using $d<p$. Since $\operatorname{Z}(\mathbf{S})\subseteq\operatorname{Z}(\Delta^{d}f)$, we have that $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(f_{d})$. [^8] Indeed, $\operatorname{Z}(\mathbf{A})=\{x\,|\,\mathbf{A}(x)=0\}=\{x\,|\,\mathbf{S}(x,\ldots,x)=0\}\subseteq\{x\,|\,\Delta^{d}f(x,\ldots,x)=0\}=\operatorname{Z}(f_{d})$.
By Corollary 3.3, $\operatorname{cor}(f_{d},P)\geq p^{-dm}$ for some polynomial $P=\sum_{i=1}^{m}c_{i}A_{i}$ of degree at most d1d-1. Note that $\operatorname{cor}(f,\,P-f_{<d})=\operatorname{bias}(f-(P-f_{<d}))=\operatorname{bias}(f_{d}-P)$. Thus,

### 30 Apr 2026 v2

@@ 2.2 Proof outline
We start by sketching the standard proof of $\operatorname{GI}_{d}(d)$. The argument is roughly as follows. For $f$ of degree $d$, the $d$-th derivative $\Delta^{d}f$ is a degree-$d$ multilinear form, which is biased by assumption. Call that bias $\delta$. Using a partition-versus-analytic rank bound, such as Theorem 2.1, one obtains a partition-rank decomposition for $\Delta^{d}f$ of length $\tilde{O}_{d}(\log_{|\mathbb{F}|}(1/\delta))$. By plugging in the diagonal $(x,\ldots,x)$ and dividing by $d!$, we obtain a decomposition of $f_{d}$ (the degree-$d$ component of $f$), meaning $\operatorname{rk}(f_{d})\leq\tilde{O}_{d}(\log_{|\mathbb{F}|}(1/\delta))$. Writing $f_{d}=\sum_{i}\alpha_{i}\beta_{i}$, a standard Fourier/pigeonhole-type argument gives a correlation for $f_{d}$ of at least $\delta^{\tilde{O}_{d}(1)}$ with some linear combination of $\alpha_{i}$ and $\beta_{i}$—which is of degree at most d1d-1—and thus a correlation for $f$ itself with such a lower-degree polynomial.
The proof of Theorem 2, for homogeneous polynomials, follows this classical recipe, except it utilizes a new correlation lemma (Lemma 3.1) to obtain a correlation with a linear combination of only the lower-degree factors from each product $\alpha_{i}\beta_{i}$ in the decomposition, which is a polynomial of degree at most $\deg(f)/2$ (<d<d).
@@
In Section 4, we introduce a new notion of rank for polynomials called $\operatorname{rk}^{*}$. The objective here was to find a notion of rank for which the following holds: whenever $\operatorname{rk}^{*}(f)$ is small, one can find a large correlation of $f$ with a polynomial of degree at most $\deg(f)-2$ (whereas for the usual notion of rank, one can only generally get $\deg(f)-1$). It turns out that the requirement needed is quite weak: for a decomposition $f=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ ($1\leq\deg(\alpha_{i})\leq\deg(\beta_{i})<\deg(f)$) to be a $\operatorname{rk}^{*}$-decomposition, we require the degree-$1$ polynomials $\alpha_{i}$ (in fact, only those with $\deg(\alpha_{i}\beta_{i})=\deg(f)$) to have linearly independent linear components. This requirement guarantees that the zero-locus of these polynomials is fairly large after an appropriate perturbation (Lemma 4.8), which is essential for us to obtain significant correlation.
We now explain how the $\operatorname{rk}^{*}$-decomposition of $f$ is obtained. Write $g=f_{d+1}$ and $h=f_{d}$ for the homogeneous components of degree $d+1$ and $d$ respectively. Proposition 5.1 provides a simple formula for $\Delta^{d}g$. Proposition 6.4 in particular shows that decomposing $g$ by the standard polarization process goes smoothly, thanks to a chain rule for bias (Lemma 6.3). The decomposition of $h$ is a more delicate issue, since the polarization $\overline{h}$ need not be biased. However, we prove that $\overline{h}$ is biased when adding an appropriate partial derivative $\partial_{c}\overline{g}$ (with respect to one of the $d$ vector variables); moreover, the sum $\overline{h}+\partial_{c}\overline{g}$ is a $d$-linear form, hence an application of Theorem 2.1 gives a decomposition to this sum. To deal with the summand $\partial_{c}\overline{g}$ we added, we use a chain rule for $\operatorname{rk}^{*}$ (Proposition 4.5) and the fact that polarization and formal differentiation commute (Lemma 5.2).
@@
We finally remark that we cannot utilize the arguments leading to the polynomial bound for $\operatorname{GI}_{d}(d)$, in Theorem 3.4, in order to obtain a similar bound for degree $d+1$, since the latter does not help in providing a decomposition, which our arguments require.
@@ 3 Correlation lemma, the homogeneous and the degree-$d$ cases
For a tuple of functions $\mathbf{A}=(A_{1},\ldots,A_{m})$, their joint zero set is $\operatorname{Z}(\mathbf{A})=\{x\,|\,\forall i\colon A_{i}(x)=0\}$. For an event $E$, we abbreviate $\operatorname{\mathfrak{bias}}(f\,|\,E)=\operatorname*{\mathbb{E}}[\chi(f)\,|\,E]$.
@@
Let $f,A_{1},\ldots,A_{m}\colon\mathbb{F}^{n}\to\mathbb{F}$ be any functions, and put $\mathbf{A}=(A_{1},\ldots,A_{m})$. Then
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\cdot\operatorname{\mathfrak{bias}}(f\,|\,\operatorname{Z}(\mathbf{A})).$
For completeness, if $\operatorname{Z}(\mathbf{A})=\emptyset$ then the right hand side should be interpreted as 0.
@@
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(f(x)-\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
$=\operatorname*{\mathbb{E}}_{x}\chi(f(x))\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\chi\Big(\sum_{i=1}^{m}c_{i}A_{i}(x)\Big)$
$=\operatorname*{\mathbb{E}}_{x}\chi(f(x))\mathbf{1}_{\operatorname{Z}(\mathbf{A})}(x)$
$=|\mathbb{F}|^{-n}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(f(x)),$
where the third equality is the fundamental fact that $\operatorname*{\mathbb{E}}_{c}\chi(c\cdot a)$ is $1$ if $a=0$, and otherwise 0 (equivalently, the orthogonality of additive characters). If $\operatorname{Z}(\mathbf{A})\neq\emptyset$, we therefore get
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)=\mathbb{P}(\operatorname{Z}(\mathbf{A}))\frac{1}{|\operatorname{Z}(\mathbf{A})|}\sum_{x\in\operatorname{Z}(\mathbf{A})}\chi(f(x))\\ =\mathbb{P}(\operatorname{Z}(\mathbf{A}))\operatorname{\mathfrak{bias}}(f\,|\,\operatorname{Z}(\mathbf{A})).\qed$
We will also need the following Theorem due to Warning (for a short proof see [Asg18]).
@@
We obtain the following correlation bound, which only makes an assumption on the zero set of $f$ (or on an ideal containing $f$). Recall the notation $\operatorname{cor}_{<d}(f)=\max_{\deg(P)<d}\operatorname{cor}(f,P)$.
@@
If $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(f)$ for some $\mathbf{A}=(A_{1},\ldots,A_{m})\in\operatorname{Form}(\mathbb{F})^{m}$ with $1\leq\deg(A_{i})<d$, then $\operatorname{cor}_{<d}(f)\geq|\mathbb{F}|^{-(d-1)m}$.
@@
Put $D=\sum_{i=1}^{m}\deg(A_{i})$. Apply Lemma 3.1 to obtain
$\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{bias}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)$ $\geq\Big|\operatorname*{\mathbb{E}}_{c\in\mathbb{F}^{m}}\operatorname{\mathfrak{bias}}\Big(f-\sum_{i=1}^{m}c_{i}A_{i}\Big)\Big|$
[…]
where the last inequality applies Warning’s second theorem (Theorem 3.2), using the fact that $0\in\operatorname{Z}(A_{1},\ldots,A_{m})$ by homogeneity. Thus, some $P=\sum_{i=1}^{m}c_{i}A_{i}\in\operatorname{Poly}_{d-1}$ satisfies $\operatorname{bias}(f-P)\geq|\mathbb{F}|^{-D}\geq|\mathbb{F}|^{-(d-1)m}$. ∎
@@
Let $f:\mathbb{F}^{n}\to\mathbb{F}$ be a homogeneous polynomial with $k:=\deg(f)<2d$, where $\mathbb{F}=\mathbb{F}_{q}$ with $\operatorname{ch}(\mathbb{F}_{q})>k$, and put $\operatorname{bias}(\Delta^{d}f)=\delta$. We assume $k\geq d\geq 2$ and $\delta>0$, as otherwise there is nothing to prove. First, we observe that $\operatorname{bias}(\overline{f})\geq\delta^{2^{d}}$, recalling that $\overline{f}=\Delta^{k}f$ denotes polarization. Indeed, using the monotonicity of Gowers norms, $\lVert{g}\rVert_{U^{k}}\geq\lVert{g}\rVert_{U^{d}}$ for $k\geq d$ and any $g\colon\mathbb{F}^{n}\to\mathbb{C}$, and using the fact that, by definition, $\operatorname{bias}(\Delta^{t}f)=\lVert{e(f)}\rVert_{U^{t}}^{2^{t}}$ for any $t$, we have that
@@
with $Q_{i},R_{i}$ multilinear forms of degrees at most k1k-1, on disjoint sets of vector variables, and
$r\leq\tilde{O}_{k}(\log_{q}(1/\operatorname{bias}(\overline{f})))=\tilde{O}_{d}(\log_{q}(1/\delta)).$
Assume without loss of generality that $\deg(Q_{i})\leq\deg(R_{i})$ for every $i$. Since $\deg(Q_{i})+\deg(R_{i})=k$, we have $\deg(Q_{i})\leq k/2$. Thus, we can put $\overline{f}$ in the ideal $\overline{f}\in\langle Q_{1},\ldots,Q_{r}\rangle$. Plug in the diagonal: let $A_{i}(x)=Q_{i}(x,\ldots,x)$, and observe that $f(x)=\frac{1}{k!}\overline{f}(x,\ldots,x)$, using the homogeneity of $f$ and $\operatorname{ch}(\mathbb{F})>k$. It follows that $f\in\langle A_{1},\ldots,A_{r}\rangle$ with each $A_{i}$ homogeneous of degree at most $k/2<d$, by assumption. We may assume $\deg(A_{i})\geq 1$, as otherwise $A_{i}=0$ so we may discard it. Since $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(f)$, we are now done by Corollary 3.3:
$\operatorname{cor}_{<d}(f)\geq q^{-dr}\geq q^{-\tilde{O}_{d}(\log_{q}(1/\delta))}=\delta^{\tilde{O}_{d}(1)}.\qed$
@@ 3.1 A remark on the case of degree-$d$ polynomials
Let $f\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(f)\leq d$ and $\operatorname{bias}(\Delta^{d}f)=\delta$, where $\mathbb{F}=\mathbb{F}_{p}$ and $d<p$. We assume $\deg(f)=d\geq 2$ and $\delta>0$, as otherwise there is nothing to prove. Recall that $\Delta^{d}f\colon(\mathbb{F}^{n})^{d}\to\mathbb{F}$ is a $d$-linear form, and let us write $\Delta^{d}f(x^{(1)},\ldots,x^{(d)})=\sum_{i=1}^{n}x^{(d)}_{i}T_{i}(x^{(1)},\ldots,x^{(d-1)})$, so that each $T_{i}\colon(\mathbb{F}^{n})^{d-1}\to\mathbb{F}$ is a (d1)(d-1)-linear form. Put $\mathbf{T}=(T_{1},\ldots,T_{n})$, and consider the multilinear variety $V=\operatorname{Z}(\mathbf{T})\subseteq(\mathbb{F}^{n})^{d-1}$. Note that the density of $V$ is $\mathbb{P}_{\mathbf{x}\in(\mathbb{F}^{n})^{d-1}}(\mathbf{x}\in V)=\operatorname{bias}(\Delta^{d}f)=\delta$; indeed, this is the standard observation $\operatorname{bias}(\Delta^{d}f)=\operatorname*{\mathbb{E}}_{\mathbf{x}=(x^{(1)},\ldots,x^{(d-1)})}\operatorname*{\mathbb{E}}_{x^{(d)}}\chi(x^{(d)}\cdot\mathbf{T}(\mathbf{x}))=\operatorname*{\mathbb{E}}_{\mathbf{x}}\mathbf{1}_{V}(\mathbf{x})=\mathbb{P}_{\mathbf{x}}(\mathbf{x}\in V)$. By Theorem 3.4 with d1d-1, there are multilinear forms $\mathbf{S}=(S_{1},\ldots,S_{m})$, with each $S_{i}\colon(\mathbb{F}^{n})^{d-1}\to\mathbb{F}$ of degrees at most d1d-1, such that $\operatorname{Z}(\mathbf{S})\subseteq\operatorname{Z}(\mathbf{T})$ and $m\leq O_{d}(\log_{p}(1/\delta)+1)$. We observe that for any nonzero $d$-linear form $T$, $\operatorname{AR}(T)\geq 2^{-d}$; this follows for example from Claim 3.3 in [Lov19]. Since $\Delta^{d}f\neq 0$, as otherwise $\deg(f)<d$, we have that $\log_{p}(1/\delta)=\operatorname{AR}(\Delta^{d}f)\geq 2^{-d}$. Thus, $\log_{p}(1/\delta)+1\leq(2^{d}+1)\log_{p}(1/\delta)$, and therefore
@@
Plugging in the diagonal, let $A_{i}(x)=S_{i}(x,\ldots,x)$, homogeneous with $\deg(A_{i})\leq\deg(S_{i})\leq d-1$. Put $\mathbf{A}=(A_{1},\ldots,A_{m})$. We have $f_{d}(x)=(1/d!)\Delta^{d}f(x,\ldots,x)$ using $d<p$. Since $\operatorname{Z}(\mathbf{S})\times\mathbb{F}^{n}\subseteq\operatorname{Z}(\mathbf{T})\times\mathbb{F}^{n}\subseteq\operatorname{Z}(\Delta^{d}f)$, we have that $\operatorname{Z}(\mathbf{A})\subseteq\operatorname{Z}(f_{d})$. [^16] Indeed, $\operatorname{Z}(\mathbf{A})=\{x\,|\,\mathbf{A}(x)=0\}=\{x\,|\,\mathbf{S}(x,\ldots,x)=0\}\subseteq\{x\,|\,\Delta^{d}f(x,\ldots,x,x)=0\}=\operatorname{Z}(f_{d})$.
We may assume $\deg(A_{i})\geq 1$, as otherwise $A_{i}=0$ so we may discard it. By Corollary 3.3, $\operatorname{cor}(f_{d},P)\geq p^{-dm}$ for some polynomial $P=\sum_{i=1}^{m}c_{i}A_{i}$ of degree at most d1d-1. Note that $\operatorname{cor}(f,\,P+f_{<d})=\operatorname{bias}(f-(P+f_{<d}))=\operatorname{bias}(f_{d}-P)$. Thus,